# Enhancing the Transferability of Adversarial Attacks on Face Recognition with Diverse Parameters Augmentation

### v1

@@ 1 Introduction
Owing to the relentless progress in deep learning, Face Recognition (FR) has achieved substantial advancements [31, 41, 7, 1, 2, 19]. Nevertheless, the susceptibility of contemporary FR models to adversarial attacks raises a significant security concern. Therefore, there is an urgent need to bolster the resilience of FR models against adversarial face examples to reveal and address the vulnerabilities.
As a result, numerous research efforts have been focused on this area. Several adversarial attacks have been developed to create adversarial face examples with features like stealthiness [30, 55, 6, 14, 32], transferability [62, 21, 63, 64], and the capacity for physical attacks [57, 56, 20]. These initiatives are aimed at improving the effectiveness of adversarial attacks on FR. Nevertheless, the transferability of these adversarial attacks remains limited. To enhance the transferability of adversarial attacks, augmentation emerges as one of the most effective methods. As illustrated in Fig. 1, augmentation-based adversarial attack consists of three augmentation types: input-based [53, 26, 47, 43], feature-based [62, 63], and parameter-based [11, 51].
@@
Technically, our proposed attack method comprises two stages: Diverse Parameters Optimization (DPO) and Hard Model Aggregation (HMA). During the DPO stage, we initiate a subset of the optimization parameters with random noise, preserving the refined parameters in the intermediate training process to yield a diverse set of surrogate models, which is instrumental in bolstering the transferability of the adversarial examples. In the HMA stage, we add beneficial perturbations [50] with optimization directions opposite to those of the adversarial perturbations onto the feature maps of the parameter-augmented surrogate models, achieving the effect of hard model augmentation [63] to further enhances the transferability of the crafted adversarial examples. Our proposed attack method effectively addresses the challenges posed by the lack of diverse parameters and the absence of the FR head. The comparison between traditional adversarial attack methods and our proposed method is illustrated in Fig. 1. By using diverse initializations for surrogate models, we can expand the parameter set of these models, thereby enhancing the transferability of the crafted adversarial examples.
@@ 2 Related Work
Adversarial Attacks on Face Recognition. Adversarial attacks on FR models can be classified based on the constraints imposed on the adversarial perturbations. These attacks are broadly categorized into two types: restricted attacks [25, 54, 8, 67, 27, 24, 4] and unrestricted attacks [5, 58, 3, 34, 37, 40, 48, 49]. Restricted attacks create adversarial examples that adhere to a specified perturbation limit, such as an $L_{p}$ norm bound. Our proposed attack method is a restricted attack. Consequently, we will delve into the specifics of restricted attacks in detail. To enhance the transferability of adversarial attacks on FR, Zhong and Deng [62] introduced DFANet, which employs dropout on the feature maps of convolutional layers to achieve an ensemble-like effect. Zhou et al. [63] proposed BPFA, enhancing attack transferability by integrating beneficial perturbations [50] onto the feature maps of FR models, resulting in the effect of hard model augmentation. Li et al. [21] leveraged additional information from FR-related tasks and applied a multi-task optimization framework to further improve the transferability of adversarial examples. Unrestricted adversarial attacks conversely generate adversarial examples without the limitations of a predefined perturbation bound. These attacks are primarily focused on physical attacks [52, 56, 20], attribute editing [30, 17], and adversarial example generation through makeup transfer [57, 14, 32, 36]. Both restricted and unrestricted adversarial attacks on FR models have substantially advanced the capabilities of these attacks. However, existing methods for generating adversarial attacks on FR often rely on surrogate models with fixed parameters, which limits the transferability of the crafted adversarial examples. In contrast, our proposed attack method addresses this limitation by diversifying the surrogate model parameters, using models with varying parameters across different epochs.
@@ 3.1 Problem Formulation and Framework
Problem Formulation. Let ${\mathcal{F}}^{vct}(\mathbf{x})\in\mathbb{R}^{r}$ denote the FR model employed by the victim to extract the embedding from a face image $\mathbf{x}$. We denote $\mathbf{x}^{s}$ and $\mathbf{x}^{t}$ as the source and target images, respectively. The objective of the adversarial attacks explored in our research is to manipulate ${\mathcal{F}}^{vct}(\mathbf{x})$ to misclassify the adversarial example $\mathbf{x}^{adv}$ as the target image $\mathbf{x}^{t}$, while ensuring that $\mathbf{x}^{adv}$ bears a close visual resemblance to $\mathbf{x}^{s}$. For clarity and brevity, the detailed optimization objective is presented in the supplementary.
In practical attack scenarios, the attacker typically cannot access the model owned by victim, denoted as ${\mathcal{F}}^{vct}$. A commonly used method to overcome this limitation is to employ a surrogate model ${\mathcal{F}}$ to generate adversarial examples and then transfer the crafted adversarial examples to the victim model [45, 63]. Consequently, the transferability becomes a critical factor in the success of the adversarial attack and constitutes the key problem studied in this research.
Framework Overview. To enhance the transferability of adversarial examples, we propose a novel method called Diverse Parameters Augmentation (DPA). DPA augments surrogate models by incorporating both pre-trained and randomly initialized parameters, resulting in a more diverse and expansive set of surrogate models. Specifically, DPA consists of two key stages: Diverse Parameters Optimization (DPO) and Hard Model Aggregation (HMA). In the DPO stage, we diversify the surrogate model parameters by combining pre-trained and random initializations, as illustrated in Fig. 2. In the HMA stage, we apply beneficial perturbations to the feature maps of the diversified surrogate models and combine them to achieve a higher degree of augmentation, as shown in Fig. 3. In the following, we will introduce our proposed DPA attack method in detail.
@@ 3.2 Diverse Parameters Optimization
Let $\mathbf{x}$ be a batch of face images, $\mathbf{w}\in\mathbb{R}^{s\times r}$ be the parameters of the FR head, $b$ be the batch size of $\mathbf{x}$, and $s$ be the class number in the training dataset. To calculate the loss function, we should first calculate the cosine similarity matrix $\mathbf{m}\in\mathbb{R}^{b\times s}$ between ${\mathcal{F}}\left(\mathbf{x}\right)$ and the parameters of the FR head $\mathbf{w}^{\top}$ using the following formula:
$\begin{gathered}\mathbf{m}=\cos{\mathbf{a}}=\widebar{\mathbf{e}}\widebar{\mathbf{w}}^{\top}\end{gathered}$ (1)
where $\widebar{\mathbf{e}}=\frac{{\mathcal{F}}\left(\mathbf{x}\right)}{\left\|{\mathcal{F}}\left(\mathbf{x}\right)\right\|}$, $\widebar{\mathbf{w}}=\frac{\mathbf{w}}{\left\|\mathbf{w}\right\|}$ and ${\mathbf{a}}$ is the angle between ${\mathcal{F}}\left(\mathbf{x}\right)$ and $\mathbf{w}^{\top}$. After obtaining the cosine similarity matrix by Eq. 1, we can get the sine similarity between ${\mathcal{F}}\left(\mathbf{x}\right)$ and $\mathbf{w}$ using the following equation:
@@
$\begin{gathered}{\mathbf{q}}=d\left(\widehat{{\mathbf{h}}}\odot{\mathbf{p}}+\left(1-\widehat{{\mathbf{h}}}\right)\odot\mathbf{m}\right)\\ {\mathbf{p}}=\left\{\begin{array}[]{ll}\cos\left({\mathbf{a}}+m\right)&\text{s.t.}\quad{\mathbf{a}}<\pi-m,\\ \cos{\mathbf{a}}-m\sin m&\text{s.t.}\quad{\mathbf{a}}\geq\pi-m,\end{array}\right.\end{gathered}$ (5)
@@
Let $c$ and ${\mathbf{v}}$ be the number of the training epoch and the parameters of the FR model, respectively. Utilizing Eq. 6, we can get the parameters set [7]:
@@
where $\varphi\left(\cdot,i\right)$ is the operation that obtain the ${\mathbf{v}}$ and ${\mathbf{w}}$ parameters in the $i$th epoch. We denote ${\mathbf{v}}_{0}$ and ${\mathbf{w}}_{0}$ as the initialized parameters of the FR model and head, respectively. If ${\mathbf{v}}_{0}$ and ${\mathbf{w}}_{0}$ are fixed, ${\mathbf{V}}_{c}$ and ${\mathbf{W}}_{c}$ become deterministic aside from minor random factors in the training process of Eq. 7. Therefore, there exists a mapping between $\left\{{\mathbf{V}}_{c},{\mathbf{W}}_{c}\right\}$ and $\left\{{\mathbf{v}}_{0},{\mathbf{w}}_{0}\right\}$ that can be expressed as:
@@
where ${\mathbf{v}}_{0}^{p}$ represents the pre-trained parameters, while ${\mathbf{w}}_{0}^{p}$, ${\mathbf{v}}_{0}^{a}$ and ${\mathbf{w}}_{0}^{a}$ are randomly initialized. Using Eq. 9, we can obtain the more diverse parameters for the FR model and head. After obtaining ${\mathbf{V}}_{c}^{p}$ and ${\mathbf{V}}_{c}^{a}$, we use the following set of the parameters to craft the adversarial examples:
@@
where $\kappa=\lfloor\sqrt{c}\rfloor$ determine the epoch interval to select the parameters. The overall framework and pseudo-code of the DPO stage is illustrated in Fig. 2 and Algorithm 1, respectively.
@@
7:           Calculate ${\mathcal{L}}$ using Eq. 6.
@@
11:        if the ${\mathbf{v}}$ satisfies the condition outlined in Eq. 10 then
@@ 3.3 Hard Model Aggregation
Crafting adversarial examples solely using vanilla models can limit the transferability. Different from the conventional method, the key innovation of the stage of our proposed attack method is the addition of beneficial perturbations [50] with the opposite optimization direction with the adversarial perturbations onto the pre-defined feature maps of the surrogate models to transform the surrogate models with diverse parameters into hard models [63] and aggregate the hard models to enhance the transferability.
@@
where $\beta$ denotes the step size used to construct the adversarial examples, while $\prod$ represents the clipping operation that confines the pixel values of the generated adversarial examples within the range $\left[\mathbf{x}^{s}-\epsilon,\mathbf{x}^{s}+\epsilon\right]$. The framework of the HMA is illustrated in Fig. 3.
@@
where $\varkappa$ represents the mapping from the pre-defined set of layers for adding beneficial perturbations $\Omega$, to their corresponding layer set. Let $f^{i}$ denote the $i$th layer within the model ${\mathcal{F}}$, and let $z$ represent the total number of layers in ${\mathcal{F}}$. We define the segment of ${\mathcal{F}}$ from layer $f^{i}$ to layer $f^{j}$ as follows:
@@ 4.1 Experimental Setting
Datasets. We opt to use the LFW [15], and CelebA-HQ [18] for our experiments to verify the effectiveness of our proposed attack method. LFW serves as an unconstrained face dataset for FR. CelebA-HQ consists of face images with high visual quality. The LFW and CelebA-HQ utilized in our experiments are identical to those employed in [63, 64, 65]. For the parameter augmentation process, we select BUPT-Balancedface [44] as our training dataset.
@@ 4.2 Comparison Studies
DPA achieves the best black-box attack results on both normally trained and adversarial robust models. To verify the effectiveness of our proposed attack method, we craft the adversarial examples on the LFW and CelebA-HQ dataset. The black-box performance with IRSE50, FaceNet, MobileFace, and IR152 as the surrogate models on the LFW dataset are demonstrated in Tab. 1, Tab. 2, Tab. 3, and Tab. 4, respectively. Some of the adversarial examples are demonstrated in Fig. 4. The results on the CelebA-HQ dataset are in the supplementary. These results demonstrate that our proposing method consistently outperforms the baseline attack, thereby highlighting its effectiveness in improving the transferability of adversarial examples.
Table 1: Comparisons of black-box ASR (%) results for attacks using IRSE50 as the surrogate model on the LFW dataset. I, S, F, M denote IR152, IRSE50, FaceNet, and MobileFace, respectively
Attacks I F M Iadv Sadv Fadv Madv
FIM 32.3 15.5 79.1 9.8 17.5 5.5 5.7
DI 59.9 47.5 97.7 25.9 41.5 15.6 23.8
DFANet 44.3 26.7 96.9 15.3 28.0 8.6 12.4
VMI 54.0 34.0 96.4 22.5 37.6 13.1 20.3
SSA 58.8 37.1 97.3 22.4 38.5 12.3 18.1
SIA 58.4 41.4 98.2 22.2 37.6 13.9 23.3
BPFA 54.4 27.5 94.6 17.6 29.4 8.1 12.8
BSR 28.7 18.4 84.5 9.2 16.3 6.5 7.6
Ours 74.4 89.8 98.3 57.9 68.2 38.3 59.7
@@
Table 2: Comparisons of black-box ASR (%) results for attacks using FaceNet as the surrogate model on the LFW dataset. I, S, F, M denote IR152, IRSE50, FaceNet, and MobileFace, respectively
Attacks I S M Iadv Sadv Fadv Madv
FIM 7.8 12.5 5.4 7.5 6.9 17.2 2.5
DI 18.6 32.2 18.5 18.0 15.8 30.2 9.9
DFANet 12.1 22.2 11.7 11.3 10.4 25.1 5.5
VMI 24.4 35.1 20.7 24.4 23.2 36.3 15.2
SSA 21.6 44.8 30.8 17.7 17.0 31.9 10.9
SIA 29.1 42.9 26.2 28.7 23.9 38.3 16.7
BPFA 17.3 31.6 14.7 13.4 13.0 22.6 7.8
BSR 28.6 42.4 25.9 26.2 24.3 34.2 16.0
Ours 42.6 65.0 56.9 47.3 45.4 54.0 31.1
Table 3: Comparisons of black-box ASR (%) results for attacks using MobileFace as the surrogate model on the LFW dataset. I, S, F, M denote IR152, IRSE50, FaceNet, and MobileFace, respectively
Attacks I S F Iadv Sadv Fadv Madv
FIM 5.3 73.4 7.5 2.5 4.5 2.8 10.9
DI 18.4 97.3 32.9 10.6 18.2 10.2 38.6
DFANet 7.0 86.4 11.9 3.6 6.4 3.9 16.8
VMI 13.6 96.0 20.2 7.6 12.8 7.6 32.3
SSA 13.8 96.4 19.6 5.5 13.1 7.2 31.5
SIA 15.7 96.8 26.7 8.1 14.5 9.1 35.3
BPFA 15.8 95.5 17.8 5.9 11.0 5.1 28.4
BSR 5.4 74.2 9.5 2.9 5.9 4.9 11.4
Ours 67.7 98.2 90.8 55.4 66.4 42.6 71.6
Table 4: Comparisons of black-box ASR (%) results for attacks using IR152 as the surrogate model on the LFW dataset. I, S, F, M denote IR152, IRSE50, FaceNet, and MobileFace, respectively
Attacks S F M Iadv Sadv Fadv Madv
FIM 29.0 9.3 5.6 13.8 6.8 3.6 2.4
DI 46.9 21.7 14.4 28.0 12.4 7.9 6.1
DFANet 50.7 15.5 12.5 25.6 11.0 5.8 3.2
VMI 49.7 23.9 18.7 30.1 18.3 12.8 11.2
SSA 55.0 21.9 24.0 28.8 14.2 9.0 6.1
SIA 52.6 26.3 19.6 29.8 18.3 11.0 9.5
BPFA 46.7 12.9 9.2 20.1 8.9 4.7 3.1
BSR 35.3 14.7 7.3 19.2 9.9 6.6 4.3
Ours 99.4 90.3 96.4 74.0 69.9 42.0 57.7
DPA achieves superior black-box performance under JPEG compression. JPEG compression is a widely adopted method for image compression during transmission, and it also serves as a defense mechanism against adversarial examples. To evaluate the effectiveness of our proposed attack under JPEG compression, we employ MobileFace as the surrogate model and IRSE50 as the victim model. We assess the attack performance on the LFW and CelebA-HQ datasets, with the results presented in Fig. 5. These results indicate that our proposed attack method consistently outperforms the baseline attack methods across varying levels of JPEG compression, thereby underscoring the robustness of our proposed attack method under such conditions.
@@
DPA demonstrates better black-box performance compared to the parameter-based augmented adversarial attack. The LGV method is an effective parameter-based augmented adversarial attack technique designed to enhance transferability. Our proposed attack incorporates parameter augmentation, making LGV an appropriate baseline. We selected MobileFace as the surrogate model and generated adversarial examples on the LFW dataset. The results are presented in Tab. 5. Tab. 5 clearly shows that our proposed attack outperforms the baseline, further validating the effectiveness of our method.
@@
DPA achieves superior black-box performance on FR models trained with various algorithms. In order to substantiate the effectiveness of our proposed attack method across various FR models, we conducted additional experiments using BSR [43] as the Baseline and FaceNet as the surrogate model. These experiments adhered to the same settings detailed in Tab. 2. The ASR across multiple FR models is illustrated in Fig. 6. As shown in Fig. 6, the ASR of the adversarial examples generated by our proposed method exceeds that of the Baseline, thereby reinforcing the efficacy of our proposed attack method.

### v3

# Improving the Transferability of Adversarial Attacks on Face Recognition with Diverse Parameters Augmentation

@@ 1 Introduction
Owing to the relentless progress in deep learning, Face Recognition (FR) has achieved substantial advancements [34, 44, 8, 1, 2, 20]. However, the inherent vulnerabilities in existing FR models pose significant security risks [7, 77, 78, 79, 31, 56, 80], with adversarial attacks emerging as one of the critical concerns. Therefore, there is an urgent need to bolster the resilience of FR models against adversarial face examples to reveal and address the vulnerabilities.
As a result, numerous research efforts have been focused on this area. Several adversarial attacks have been developed to create adversarial face examples with features like stealthiness [33, 64, 6, 15, 35], transferability [72, 22, 74, 73], and the capacity for physical attacks [66, 65, 21]. These initiatives are aimed at improving the effectiveness of adversarial attacks on FR. Nevertheless, the transferability of these adversarial attacks remains limited. To enhance the transferability of adversarial attacks, augmentation emerges as one of the most effective methods. As illustrated in Figure 1, augmentation-based adversarial attack consists of three augmentation types: input-based [62, 27, 55, 46], feature-based [72, 74], and parameter-based [12, 60].
@@
Technically, our proposed attack method comprises two stages: Diverse Parameters Optimization (DPO) and Hard Model Aggregation (HMA). During the DPO stage, we initiate a subset of the optimization parameters with random noise, preserving the refined parameters in the intermediate training process to yield a diverse set of surrogate models, which is instrumental in bolstering the transferability of the adversarial examples. In the HMA stage, we add beneficial perturbations [59] with optimization directions opposite to those of the adversarial perturbations onto the feature maps of the parameter-augmented surrogate models, achieving the effect of hard model augmentation [74] to further enhances the transferability of the crafted adversarial examples. Our proposed attack method effectively addresses the challenges posed by the lack of diverse parameters and the absence of the FR head. The comparison between traditional adversarial attack methods and our proposed method is illustrated in Figure 1. By using diverse initializations for surrogate models, we can expand the parameter set of these models, thereby enhancing the transferability of the adversarial examples.
@@ 2 Related Work
Adversarial Attacks on Face Recognition. Adversarial attacks on FR models can be classified based on the constraints imposed on the adversarial perturbations. These attacks are broadly categorized into two types: restricted attacks [26, 63, 9, 82, 28, 25, 4, 75] and unrestricted attacks [5, 67, 3, 37, 40, 43, 57, 58]. Restricted attacks create adversarial examples that adhere to a specified perturbation limit, such as an $L_{p}$ norm bound. Our proposed attack method is a restricted attack. Consequently, we will delve into the specifics of restricted attacks in detail. To enhance the transferability of adversarial attacks on FR, Zhong and Deng [72] introduced DFANet, which employs dropout on the feature maps of convolutional layers to achieve an ensemble-like effect. Zhou et al. [74] proposed BPFA, enhancing attack transferability by integrating beneficial perturbations [59] onto the feature maps of FR models, resulting in the effect of hard model augmentation. Li et al. [22] leveraged additional information from FR-related tasks and applied a multi-task optimization framework to further improve the transferability of adversarial examples. Unrestricted adversarial attacks conversely generate adversarial examples without the limitations of a predefined perturbation bound. These attacks are primarily focused on physical attacks [61, 65, 21], attribute editing [33, 18], and adversarial example generation through makeup transfer [66, 15, 35, 39]. Both restricted and unrestricted adversarial attacks on FR models have substantially advanced the capabilities of these attacks. However, existing methods for generating adversarial attacks on FR often rely on surrogate models with fixed parameters, which limits the transferability of the crafted adversarial examples. In contrast, our proposed attack method addresses this limitation by diversifying the surrogate model parameters, using models with varying parameters across different epochs.
@@ 3.1 Problem Formulation and Framework
Problem Formulation. Let ${\mathcal{F}}^{vct}(\mathbf{x}^{\prime})\in\mathbb{R}^{r}$ denote the FR model employed by the victim to extract the embedding from a face image $\mathbf{x}^{\prime}$. We denote $\mathbf{x}^{s}$ and $\mathbf{x}^{t}$ as the source and target images, respectively. The objective of the adversarial attacks explored in our research is to manipulate ${\mathcal{F}}^{vct}(\mathbf{x})$ to misclassify the adversarial example $\mathbf{x}^{adv}$ as the target image $\mathbf{x}^{t}$, while ensuring that $\mathbf{x}^{adv}$ bears a close visual resemblance to $\mathbf{x}^{s}$. The objective of adversarial attack as delineated in this paper is as follows:
In practice, the attacker typically cannot access the model owned by the victim. As a result, black-box attacks play a crucial role in adversarial attacks [51, 52, 32]. A widely adopted approach for black-box attacks involves using a surrogate model ${\mathcal{F}}$ to generate adversarial examples and then transfer the crafted adversarial examples to the victim model [53, 74]. Consequently, the transferability is important and constitutes the key problem studied in this research.
Framework Overview. To enhance the transferability of adversarial examples, we propose a novel method called Diverse Parameters Augmentation (DPA). DPA augments surrogate models by incorporating both pre-trained and randomly initialized parameters, resulting in a more diverse and expansive set of surrogate models. Specifically, DPA consists of two key stages: Diverse Parameters Optimization (DPO) and Hard Model Aggregation (HMA). In the DPO stage, we diversify the surrogate model parameters by combining pre-trained and random initializations, as illustrated in Figure 2. In the HMA stage, we apply beneficial perturbations to the feature maps of the diversified surrogate models and combine them to achieve a higher degree of augmentation, as shown in Figure 3. In the following, we will introduce our proposed DPA attack method in detail.
@@ 3.2 Diverse Parameters Optimization
Let $\mathbf{x}$ be a batch of face images, $\mathbf{w}\in\mathbb{R}^{s\times r}$ be the parameters of the FR head, $b$ be the batch size of $\mathbf{x}$, and $s$ be the class number in the training dataset. To calculate the loss function, we should first calculate the cosine similarity matrix ${\mathbf{r}}\in\mathbb{R}^{b\times s}$ between ${\mathcal{F}}\left(\mathbf{x}\right)\in\mathbb{R}^{b\times r}$ and the parameters of the FR head $\mathbf{w}^{\top}$ using the following formula:
$\begin{gathered}{\mathbf{r}}=\cos{\mathbf{a}}=\widebar{\mathbf{e}}\widebar{\mathbf{w}}^{\top}\end{gathered}$ (2)
where $\widebar{\mathbf{e}}=\frac{{\mathcal{F}}\left(\mathbf{x}\right)}{\left\|{\mathcal{F}}\left(\mathbf{x}\right)\right\|}$, $\widebar{\mathbf{w}}=\frac{\mathbf{w}}{\left\|\mathbf{w}\right\|}$ and ${\mathbf{a}}$ is the angle between ${\mathcal{F}}\left(\mathbf{x}\right)$ and $\mathbf{w}^{\top}$. After obtaining the cosine similarity matrix by Equation 2, we can get the sine similarity between ${\mathcal{F}}\left(\mathbf{x}\right)$ and $\mathbf{w}$ using the following equation:
@@
$\begin{gathered}{\mathbf{q}}=d\left(\widehat{{\mathbf{h}}}\odot{\mathbf{p}}+\left(1-\widehat{{\mathbf{h}}}\right)\odot{\mathbf{r}}\right)\\ {\mathbf{p}}=\left\{\begin{array}[]{ll}\cos\left({\mathbf{a}}+m\right)&\text{s.t.}\quad{\mathbf{a}}<\pi-m,\\ \cos{\mathbf{a}}-m\sin m&\text{s.t.}\quad{\mathbf{a}}\geq\pi-m,\end{array}\right.\end{gathered}$ (6)
@@
Let $c$ and ${\mathbf{v}}$ be the number of the training epoch and the parameters of the FR backbone, respectively. Utilizing Equation 7, we can get the parameters set [8]:
@@
where $\varphi\left(\cdot,i\right)$ is the operation that obtain the ${\mathbf{v}}$ and ${\mathbf{w}}$ parameters in the $i$th epoch. We denote ${\mathbf{v}}_{0}$ and ${\mathbf{w}}_{0}$ as the initialized parameters of the FR model and head, respectively. If ${\mathbf{v}}_{0}$ and ${\mathbf{w}}_{0}$ are fixed, ${\mathbf{V}}_{c}$ and ${\mathbf{W}}_{c}$ become deterministic aside from minor random factors in the training process of Equation 8. Therefore, there exists a mapping between $\left\{{\mathbf{V}}_{c},{\mathbf{W}}_{c}\right\}$ and $\left\{{\mathbf{v}}_{0},{\mathbf{w}}_{0}\right\}$ that can be expressed as:
@@
where ${\mathbf{v}}_{0}^{p}$ represents the pre-trained parameters, while ${\mathbf{w}}_{0}^{p}$, ${\mathbf{v}}_{0}^{a}$ and ${\mathbf{w}}_{0}^{a}$ are randomly initialized. Using Equation 10, we can obtain the more diverse parameters for the FR model and head. After obtaining ${\mathbf{V}}_{c}^{p}$ and ${\mathbf{V}}_{c}^{a}$, we use the following set of the parameters to craft the adversarial examples:
@@
where $\kappa=\lfloor\sqrt{c}\rfloor$ determines the epoch interval to select the parameters. The overall framework and pseudo-code of the DPO stage in our proposed method is illustrated in Figure 2 and Algorithm 1, respectively.
@@
7:           Calculate ${\mathcal{L}}$ using Equation 7.
@@
11:        if the ${\mathbf{v}}$ satisfies the condition outlined in Equation 11 then
@@ 3.3 Hard Model Aggregation
Hard models [74] refer to models with higher loss values compared to normal models, designed to simulate the effect of hard samples, which are hard to the optimization process of the adversarial examples. Since hard samples can improve the generalization of models, we leverage multiple hard models to enhance the transferability of the surrogate models with diverse parameters by adding beneficial perturbations [59] onto the pre-defined feature maps of the surrogate models.
@@
where $\beta$ denotes the step size used to construct the adversarial examples, while $\prod$ represents the clipping operation that confines the pixel values of the generated adversarial examples within the range $\left[\mathbf{x}^{s}-\epsilon,\mathbf{x}^{s}+\epsilon\right]$. The framework of the HMA is illustrated in Figure 3.
@@
where $\varkappa$ represents the mapping from the pre-defined set of layers for adding beneficial perturbations $\Omega$, to their corresponding layer index set $\Phi$. Let $f^{i}$ denote the $i$th layer within the model ${\mathcal{F}}$, and let $z$ represent the total number of layers in ${\mathcal{F}}$. We define the segment of ${\mathcal{F}}$ from layer $f^{i}$ to layer $f^{j}$ as follows:
@@ 4.1 Experimental Setting
Datasets. We opt to use the LFW [16], and CelebA-HQ [19] for our experiments to verify the effectiveness of our proposed attack method. LFW serves as an unconstrained face dataset for FR. CelebA-HQ consists of face images with high visual quality. The LFW and CelebA-HQ utilized in our experiments are identical to those employed in [74, 73, 76]. For the parameter augmentation process, we select BUPT-Balancedface [49, 50, 47, 48] as our training dataset.
@@ 4.2 Comparison Studies
DPA achieves the best black-box attack results on both normally trained and adversarial robust models. To verify the effectiveness of our proposed attack method, we craft the adversarial examples on the LFW and CelebA-HQ dataset. The black-box performance with MobileFace as the surrogate models on the LFW dataset are demonstrated in Table 1. Some of the adversarial examples are demonstrated in Figure 4. More experimental results on the LFW dataset, as well as those on the CelebA-HQ dataset, are provided in the supplementary. These results demonstrate that our proposing method consistently outperforms the baseline attack, thereby highlighting its effectiveness in improving the transferability of adversarial face examples.
@@
Table 1: Comparisons of black-box ASR (%) results for attacks using MobileFace as the surrogate model on the LFW dataset. I, S, F, M denote IR152, IRSE50, FaceNet, and MobileFace, respectively.
Attacks I S F Iadv Sadv Fadv Madv FIM [71] 5.3 73.4 7.5 2.5 4.5 2.8 10.9 DI [62] 18.4 97.3 32.9 10.6 18.2 10.2 38.6 DFANet [72] 7.0 86.4 11.9 3.6 6.4 3.9 16.8 VMI [53] 13.6 96.0 20.2 7.6 12.8 7.6 32.3 SSA [27] 13.8 96.4 19.6 5.5 13.1 7.2 31.5 SIA [55] 15.7 96.8 26.7 8.1 14.5 9.1 35.3 BPFA [74] 15.8 95.5 17.8 5.9 11.0 5.1 28.4 BSR [46] 5.4 74.2 9.5 2.9 5.9 4.9 11.4 Ours 67.7 98.2 90.8 55.4 66.4 42.6 71.6
DPA achieves superior black-box performance under JPEG compression. JPEG compression is a widely adopted method for image compression during transmission, and it also serves as a defense mechanism against adversarial examples. To evaluate the effectiveness of our proposed attack under JPEG compression, we employ MobileFace as the surrogate model and IRSE50 as the victim model. We assess the attack performance on the LFW and CelebA-HQ datasets, with the results presented in Figure 5. These results indicate that our proposed attack method consistently outperforms the baseline attack methods across varying levels of JPEG compression, thereby underscoring the robustness of our proposed attack method under such conditions.
@@
DPA demonstrates better black-box performance compared to the parameter-based augmented adversarial attack. The LGV method is an effective parameter-based augmented adversarial attack technique designed to enhance transferability. Our proposed attack incorporates parameter augmentation, making LGV an appropriate baseline. We selected MobileFace as the surrogate model and generated adversarial examples on the LFW dataset. The results are presented in Table 2. Table 2 clearly shows that our proposed attack outperforms the baseline, further validating the effectiveness of our method. The primary limitation of LGV stems from its random selection of one model per iteration, which introduces instability and adversely impacts black-box ASR [9, 53]. Furthermore, LGV suffers from a lack of diverse initialization parameters. These shortcomings significantly impair the overall ASR performance of LGV.
@@
DPA achieves superior black-box performance on FR models trained with various algorithms. In order to substantiate the effectiveness of our proposed attack method across various FR models, we conducted additional experiments using BSR [46] as the Baseline and FaceNet as the surrogate model. These experiments adhered to the same settings detailed in Table 7 in the supplementary. The ASR across multiple FR models is illustrated in Figure 6. As shown in Figure 6, the ASR of the adversarial examples generated by our proposed method exceeds that of the Baseline, thereby reinforcing the efficacy of our proposed method.